OpenStudy (anonymous):

@mathsciencehistory 2+2-3+(-14)

OpenStudy (anonymous):

.

OpenStudy (unanimoose):

+(-14+ is the same as just -14.

OpenStudy (anonymous):

hi

OpenStudy (unanimoose):

2 + 2 - 3 -14

OpenStudy (anonymous):

2 + 2 = 4 - 3 = 1 + (-14) = ?

OpenStudy (anonymous):

-15?

OpenStudy (anonymous):

nope. try again, plz !

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

aren't you adding negatives?

OpenStudy (unanimoose):

1 - 14

OpenStudy (anonymous):

you can;t give answers, it is against the rules., @flooder .

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

nobody care about these rules i keep hearing of

OpenStudy (unanimoose):

?

OpenStudy (anonymous):

well @mathsciencehistory , we are tying to help them, not give them the answers so they don't know how to get it the next time.

OpenStudy (anonymous):

-3+(-14)=-3-14=-17 What is 4-17 ?

OpenStudy (anonymous):

2 + 2 = 4 - 3 = 1 + (-14) = ? basically, it is down to 1 + (-14) 1 + (-14) can turn into 1 - 14 1 - 14 = ?

OpenStudy (anonymous):

+(-14) is -14 -(+14) is -14 minus changes from - to + or + to -, but + does not have any effect